OpenStudy (anonymous):

i need the innequality 3x + 5 is less than or equal to 2x - 8

OpenStudy (anonymous):

3x + 5 <= 2x - 8 (subtract 2x from both sides) 3x - 2x + 5 <= -8 (subtract 5 from both sides) 3x - 2x <= -8 - 5 (combine like terms) x <= - 13